OpenStudy (anonymous):

6y=6 would i -6 into the other 6 to make y=0 or am i wrong and how so i fix it

OpenStudy (ja1):

How would you isolate the variable here?

OpenStudy (anonymous):

minute 6 on both sides

OpenStudy (ja1):

Not quite. Remember that \[6y\] is also the same as \[6(y)\]Which is multiplication.

OpenStudy (ja1):

Now you need to do the inverse operation. What is the inverse of multiplication?

OpenStudy (anonymous):

No you are supposed to the opposite of what is originally done. In that equation it is multiplying so you would divide 6 by 6.

OpenStudy (anonymous):

ok i get the multiplying so i would write 6times 6 then

OpenStudy (ja1):

No. Because 6y is already you need to do the inverse (opposite) which is, like Kiara said, division. So you need to divide 6y by 6. Can you do that for me?

OpenStudy (ja1):

already multiplication* sorry

OpenStudy (anonymous):

so 6/6 that would equal 1 and then the answer would be y=1

OpenStudy (ja1):

Yes! Good job.

OpenStudy (anonymous):

ok thank you for all ur help i seem to better understand how to do these math problem with people like u helping ,e because i dont get much help at school thats why i bring my homwork home and get help online.

OpenStudy (ja1):

That's what we're here for! Glad I could help :)

OpenStudy (anonymous):

now im struggling with finding the number n by writing an solving and equation.

OpenStudy (anonymous):

Eight more than a number is fifteen. i need help in figuring out to find the number

OpenStudy (ja1):

This can be expressed as follows: 8 + x = 15

OpenStudy (ja1):

Now what is the inverse of 8 + x ?

OpenStudy (anonymous):

i would say 8- 15 i think

OpenStudy (ja1):

Yes, the inverse is subtraction so now divide 8 on both sides.

OpenStudy (anonymous):

which would equal 7 if i did it right

OpenStudy (ja1):

Yes! now let's check our work: 7 + 8 = 15

OpenStudy (ja1):

So it's right :)

OpenStudy (anonymous):

-8 into 15 then 8+7+ 15 thats how i have it wrote out

OpenStudy (ja1):

However you do it, just as long as it's right :)

OpenStudy (anonymous):

could you possibly help me try another way of writing it to where i understand it better then how i wrote it

OpenStudy (anonymous):

i wrote it out like this n-8=15 then i got stuck from there